# Assessing the Causal Impact of Humanitarian Aid on Food Security

### v2

@@ Abstract
In the face of climate change-induced droughts, vulnerable regions encounter severe threats to food security, demanding urgent humanitarian assistance. This paper introduces a causal inference framework for the Horn of Africa, aiming to assess the impact of cash-based interventions on food crises. Our contributions include identifying causal relationships within the food security system, harmonizing a comprehensive database including socio-economic, weather and remote sensing data, and estimating the causal effect of humanitarian interventions on malnutrition. On a country level, our results revealed no significant effects, likely due to limited sample size, suboptimal data quality, and an imperfect causal graph resulting from our limited understanding of multidisciplinary systems like food security. Instead, on a district level, results revealed significant effects, further implying the context-specific nature of the system. This underscores the need to enhance data collection and refine causal models with domain experts for more effective future interventions and policies, improving transparency and accountability in humanitarian aid.
@@ 1 Introduction
In a world where climate change is rapidly accelerating, droughts are becoming more frequent and severe, posing a serious challenge to food security in the most vulnerable regions of our planet. In this context, communities that rely solely on rainfall for their livelihoods are especially at risk, often requiring immediate humanitarian assistance to survive [1, 2]. Failure to act or provide adequate aid can have immense consequences, including devastating economic losses, mass displacement of people, malnutrition in infants, and elevated mortality rates due to hunger and famine [3, 4, 5]. Humanitarian organizations are facing a significant challenge due to the widening gap between funding and the needs of the people affected by food crises [6, 7]. As a result, designing effective humanitarian interventions in resource-constrained situations has become a critical issue. Despite numerous comprehensive reviews, there is still a lack of solid evidence to identify the best strategies to help populations affected by crises [8]. Cash-based and voucher aid programs are considered effective in emergencies, but their cost-effectiveness varies by context [9]. Standardized methods for evaluating humanitarian interventions in food emergencies are lacking [8]. Our aim is to determine the impact of interventions, using observational causal inference to enhance intervention design, and transparency in charity, and improve humanitarian aid outcomes during extreme droughts.
The Horn of Africa has witnessed a concerning rise in acute malnutrition, affecting 6.5 million people in 2022 [6]. Prolonged dry spells significantly contribute to this crisis [10], yet it is crucial to recognize that droughts are not the sole driver. Various factors, including hydrological conditions, food production capabilities, market access, insufficient humanitarian aid, conflicts, and displacement, play a significant role [11, 12, 13, 14, 15]. Studying food security in this context is intricate, involving multiple variables, scales, and non-linear relationships. Predictive Machine Learning (ML) techniques are not suited to understanding the causes and estimating the causal effect by default [16, 17], instead, this paper focuses on causal inference, specifically assessing the impact of humanitarian interventions during the 2016, 2018, and 2022 Horn of Africa droughts. Our aim is to demonstrate the application of causal inference for evaluating the effectiveness of cash-based interventions in food crisis scenarios.
@@ 2 Related Work
In recent years, the surge in available data has enabled us to assess the impact of climate change on food insecurity. This data originates from diverse sources, encompassing Earth observation products [18] and systematic socioeconomic data collection programs [19, 20]. Leveraging this wealth of data, we can estimate causal effects from observations [16, 17]. This approach is particularly vital in domains where conducting controlled experiments is impractical, costly, or unethical, with food insecurity research being a prominent example. Observational data for causal inference has gained prominence across various disciplines, including ecology [21], agriculture [22, 23, 24], public policy [25, 26], and Earth sciences [27, 28]. While there have been subjective and technical assessments of humanitarian interventions in emergency contexts [8, 29], to the best of our knowledge, this is the first effort to apply modern observational causal inference methods to evaluate humanitarian policy in a food emergency context. It is also the first time such a broad database of driving factors has been used for this purpose. The contributions of our work are summarized as follows: i) identifying the overarching causal graph and the drivers of food insecurity in the Horn of Africa, ii) building a harmonized database with the best available data suitable to evaluate cash-based interventions, iii) the estimation of the causal effect of humanitarian interventions on malnutrition.
@@ 3 Causal Inference
To reason about the causal effects of certain random variables on others, first, we need to codify causal relations. Causal inference provides a language for formalizing structural knowledge about the data-generating process [37] with which we can estimate what will happen to data after changes, called interventions. The canonical representation of causal relations is a causal Direct Acyclic Graph (DAG), which can encode a priori assumptions about the causal structure of interest. In causal modeling, assumptions are crucial, as establishing relations solely on observational data can prove challenging. Background knowledge and domain expertise are a common source of assumptions in causal inference. Randomized controlled trials (RCTs) are also a gold standard for establishing causality because, under certain conditions, random assignment helps control for confounding variables. However, RCTs also bring to attention when this sort of experiment cannot be performed. When studying complex dynamic systems such as the food security system, replicating interventional experiments could prove infeasible or unethical. Therefore, when modeling the causal relations of the food security system involving climate and socio-economic dynamics, multi-scale and non-linear drivers, we rely solely on the information provided by background knowledge and associated literature.
@@ 4 Data and Methods
Fig. 1: DAG representing the food security system in Somalia.
@@ 4.2 Data
Food security is influenced by various climatic, economic, and social factors, as represented in our DAG, which reflects the dynamics of agropastoralist households in drought displacement situations in Somalia [42]. We collect and harmonize data for the variables in the DAG from multiple sources (Table 1). Our outcome variable is the food security index GAM [19], and the treatment variable is a proxy for cash interventions, reflecting the number of individuals who received money in the form of credit or remittances [19]. We also collect data on El Niño Southern Oscillation (ENSO) to account for climate variability [43] and use the Standardized Precipitation Index (SPI) to characterize dry spells [44, 18]. Socio-economic data include monthly market prices of livestock, staple food, water, and sorghum production [45]. We measure conflict levels using a proxy based on recorded fatalities [46] and incorporate data on drought-induced internal displacement [20]. All data are aggregated annually and by district.
@@ 4.3 Problem Formulation
To estimate the Average Treatment Effect (ATE), $\text{ATE}=\mathop{\mathbb{E}}[Y|do(T=1)]-\mathop{\mathbb{E}}[Y|do(T=0)]$, we identify an adjustment set $Z\subseteq V$. We apply the back-door criterion, which relies on a graphical test to determine whether adjusting for a set of graph nodes $Z\subseteq V$ is sufficient for estimating $\mathbb{P}(Y=y|do(T=t))$. We find the parent adjustment set that is sufficient for estimating the ATE, which is $Z=${Market Prices, Sorghum Production, Fatalities, Drought-induced internal displacements, Population}. Utilizing the Potential Outcomes framework, our ATE estimation aims to capture the difference between the average GAM values under humanitarian aid exceeding a chosen threshold and the average value of the outcome when humanitarian aid falls below that threshold. To estimate the effect, we use several methods of varying complexity. Linear regression (LR) and distance matching (M) are selected as baseline estimation methods. The popular Inverse Propensity Score weighting (IPS W) is also used [47], as well as modern machine learning methods, the T-learner (T-L) and X-learner (X-L) [48]. Given the unavailability of observed ground truth estimates, we resort to performing refutation tests, in line with recent research [49, 50], to assess the robustness of our models. We perform the following tests: i) Placebo treatment, where the treatment is randomly permuted, and the estimated effect is expected to drop to 0; ii) Random Common Cause (RCC), where a random confounder is added to the dataset and the estimate is expected to remain unchanged; iii) Random Subset Removal (RSR), where a subset of data is randomly removed and the effect is expected to remain the same.
@@ 6 CauseMe platform
Given the data quality challenges and the context-driven nature of food security systems, our focus turns to data-driven causal discovery as a crucial methodology. The complexity of multi-domain issues, such as food security, renders traditional expertise insufficient for constructing robust DAGs. Enter the innovative CauseMe platform [52], functioning as the link between causal inference experts and domain specialists. This platform empowers non-causality experts to employ data-driven causal discovery methods, facilitating data exploration and initial DAG construction. Developed as an accessible tool for causal discovery methods, the platform [53] aims to democratize access to these techniques across diverse scientific fields. Catering to experts less versed in causality but eager to conduct further data experiments, CauseMe allows the execution of various causal discovery methods on time series data through an interactive interface. Users can tweak method parameters, select variables for causal analysis, and obtain graphical representations showcasing the returned causal relationships. Moreover, CauseMe facilitates the interpretation of these graphs through a Large Language Model (LLM), allowing users to input contextual information about the data, thereby receiving explanations of the obtained graph.
@@ 7 Conclusions
Optimally distributing available resources and evaluating how, who, where, and when to intervene is crucial to mitigating climate change impacts. In this paper, we presented a novel data-driven approach for assessing the effectiveness of humanitarian interventions in food emergencies through the lens of causal inference. We constructed a DAG to capture the dynamics of food insecurity under drought conditions and collected data characterizing the system. Our goal was to estimate the causal effects of cash-based interventions on reducing district-level food insecurity across Somalia. Preliminary country-wise results did not reach statistical significance, although a singular district analysis did, prompting further steps: i) identifying more suitable treatment variables, ii) refining the causal graph with domain experts, iii) gaining insights on the spatio-temporal heterogeneity of impact of interventions through Conditional Average Treatment Effects (CATE) [54]. If data allows it, causal inference can be used to assess the efficacy of interventions in specific locations, supporting targeted aid where on-ground surveys are not feasible. The proposed approach could promote greater accountability and transparency amongst humanitarian actors, encouraging individuals to contribute to impactful and traceable aid.
@@ 8 Acknowledgments
This work was supported by the Fundación BBVA with the project ‘Causal inference in the human biosphere coupled system (SCALE)’ and the Microsoft Climate Research Initiative through the Causal4Africa project.

### v3

@@ Abstract
In the face of climate change-induced droughts, vulnerable regions encounter severe threats to food security, demanding urgent humanitarian assistance. This paper introduces a causal inference framework for the Horn of Africa, aiming to assess the impact of cash-based interventions on food crises. Our contributions include identifying causal relationships within the malnutrition system, harmonizing a comprehensive database including socio-economic, weather and remote sensing data, and estimating the causal effect of cash-based interventions on malnutrition. On a country level, our results revealed no significant effects, likely due to limited sample size, suboptimal data quality, and an imperfect causal graph resulting from our limited understanding of multidisciplinary systems like malnutrition. Instead, on a district level, results revealed significant effects, further implying the context-specific nature of the system. This underscores the need to enhance data collection and refine causal models with domain experts for more effective future interventions and policies, improving transparency and accountability in humanitarian aid.
@@ 1 Introduction
In a world where climate change is rapidly accelerating, droughts are becoming more frequent and severe, posing a serious challenge to food security in the most vulnerable regions of our planet. In this context, communities that rely solely on rainfall for their livelihoods are especially at risk, often requiring immediate humanitarian assistance to survive [1, 2]. Failure to act or provide adequate aid can have immense consequences, including devastating economic losses, mass displacement of people, malnutrition in infants, and elevated mortality rates due to hunger and famine [3, 4, 5]. Humanitarian organizations are facing a significant challenge due to the widening gap between funding and the needs of the people affected by food crises [6, 7]. As a result, designing effective humanitarian interventions in resource-constrained situations has become a critical issue. Despite numerous comprehensive reviews, there is still a lack of solid evidence to identify the best strategies to help populations affected by crises [8]. Cash-based and voucher aid programs are considered effective in emergencies, but their cost-effectiveness varies by context [9]. Standardized methods for evaluating cash-based interventions in food emergencies are lacking [8]. Our aim is to determine the impact of interventions, using observational causal inference to enhance intervention design, and transparency in charity, and improve cash-based aid outcomes during extreme droughts.
The Horn of Africa has witnessed a concerning rise in acute malnutrition, affecting 6.5 million people in 2022 [6]. Prolonged dry spells significantly contribute to this crisis [10], yet it is crucial to recognize that droughts are not the sole driver. Various factors, including hydrological conditions, food production capabilities, market access, insufficient aid, conflicts, and displacement, play a significant role [11, 12, 13, 14, 15]. Studying malnutrition in this context is intricate, involving multiple variables, scales, and non-linear relationships. Predictive Machine Learning (ML) techniques are not suited to understanding the causes and estimating the causal effect by default [16, 17], instead, this paper focuses on causal inference, specifically assessing the impact of cash-based interventions during the 2016, 2018, and 2022 Horn of Africa droughts. Our aim is to demonstrate the application of causal inference for evaluating the effectiveness of cash-based interventions in malnutrition scenarios.
@@ 2 Related Work
In recent years, the surge in available data has enabled us to assess the impact of climate change on food insecurity. This data originates from diverse sources, encompassing Earth observation products [18] and systematic socioeconomic data collection programs [19, 20]. Leveraging this wealth of data, we can estimate causal effects from observations [16, 17]. This approach is particularly vital in domains where conducting controlled experiments is impractical, costly, or unethical, with food insecurity research being a prominent example. Observational data for causal inference has gained prominence across various disciplines, including ecology [21], agriculture [22, 23, 24], public policy [25, 26], and Earth sciences [27, 28]. While there have been subjective and technical assessments of cash-based interventions in emergency contexts [8, 29], to the best of our knowledge, this is the first effort to apply modern observational causal inference methods to evaluate the effectiveness of cash-based aid in a food emergency context. It is also the first time such a broad database of driving factors has been used for this purpose. The contributions of our work are summarized as follows: i) identifying the overarching causal graph and the drivers of malnutrition in the Horn of Africa, ii) building a harmonized database with the best available data suitable to evaluate cash-based interventions, iii) the estimation of the causal effect of cash-based interventions on malnutrition.
@@ 3 Causal Inference
To reason about the causal effects of certain random variables on others, first, we need to codify causal relations. Causal inference provides a language for formalizing structural knowledge about the data-generating process [37] with which we can estimate what will happen to data after changes, called interventions. The canonical representation of causal relations is a causal Direct Acyclic Graph (DAG), which can encode a priori assumptions about the causal structure of interest. In causal modeling, assumptions are crucial, as establishing relations solely on observational data can prove challenging. Background knowledge and domain expertise are a common source of assumptions in causal inference. Randomized controlled trials (RCTs) are also a gold standard for establishing causality because, under certain conditions, random assignment helps control for confounding variables. However, RCTs also bring to attention when this sort of experiment cannot be performed. When studying complex dynamic systems such as the malnutrition system, replicating interventional experiments could prove infeasible or unethical. Therefore, when modeling the causal relations of the malnutrition system involving climate and socio-economic dynamics, multi-scale and non-linear drivers, we rely solely on the information provided by background knowledge and associated literature.
@@ 4 Data and Methods
Fig. 1: DAG representing the malnutrition system in Somalia.
@@ 4.2 Data
Malnutrition is influenced by various climatic, economic, and social factors, as represented in our DAG, which reflects the dynamics of agropastoralist households in drought displacement situations in Somalia [42]. We collect and harmonize data for the variables in the DAG from multiple sources (Table 1). Our outcome variable is the malnutrition index GAM [19], and the treatment variable is a proxy for cash interventions, reflecting the number of individuals who received money in the form of credit or remittances [19]. We also collect data on El Niño Southern Oscillation (ENSO) to account for climate variability [43] and use the Standardized Precipitation Index (SPI) to characterize dry spells [44, 18]. Socio-economic data include monthly market prices of livestock, staple food, water, and sorghum production [45]. We measure conflict levels using a proxy based on recorded fatalities [46] and incorporate data on drought-induced internal displacement [20]. All data are aggregated annually and by district.
@@ 4.3 Problem Formulation
To estimate the Average Treatment Effect (ATE), $\text{ATE}=\mathop{\mathbb{E}}[Y|do(T=1)]-\mathop{\mathbb{E}}[Y|do(T=0)]$, we identify an adjustment set $Z\subseteq V$. We apply the back-door criterion, which relies on a graphical test to determine whether adjusting for a set of graph nodes $Z\subseteq V$ is sufficient for estimating $\mathbb{P}(Y=y|do(T=t))$. We find the parent adjustment set that is sufficient for estimating the ATE, which is $Z=${Market Prices, Sorghum Production, Fatalities, Drought-induced internal displacements, Population}. Utilizing the Potential Outcomes framework, our ATE estimation aims to capture the difference between the average GAM values under cash-based aid exceeding a chosen threshold and the average value of the outcome when cash-based aid falls below that threshold. To estimate the effect, we use several methods of varying complexity. Linear regression (LR) and distance matching (M) are selected as baseline estimation methods. The popular Inverse Propensity Score weighting (IPS W) is also used [47], as well as modern machine learning methods, the T-learner (T-L) and X-learner (X-L) [48]. Given the unavailability of observed ground truth estimates, we resort to performing refutation tests, in line with recent research [49, 50], to assess the robustness of our models. We perform the following tests: i) Placebo treatment, where the treatment is randomly permuted, and the estimated effect is expected to drop to 0; ii) Random Common Cause (RCC), where a random confounder is added to the dataset and the estimate is expected to remain unchanged; iii) Random Subset Removal (RSR), where a subset of data is randomly removed and the effect is expected to remain the same.
@@ 6 CauseMe platform
Given the data quality challenges and the context-driven nature of malnutrition systems, our focus turns to data-driven causal discovery as a crucial methodology. The complexity of multi-domain issues, such as malnutrition, renders traditional expertise insufficient for constructing robust DAGs. Enter the innovative CauseMe platform [52], functioning as the link between causal inference experts and domain specialists. This platform empowers non-causality experts to employ data-driven causal discovery methods, facilitating data exploration and initial DAG construction. Developed as an accessible tool for causal discovery methods, the platform [53] aims to democratize access to these techniques across diverse scientific fields. Catering to experts less versed in causality but eager to conduct further data experiments, CauseMe allows the execution of various causal discovery methods on time series data through an interactive interface. Users can tweak method parameters, select variables for causal analysis, and obtain graphical representations showcasing the returned causal relationships. Moreover, CauseMe facilitates the interpretation of these graphs through a Large Language Model (LLM), allowing users to input contextual information about the data, thereby receiving explanations of the obtained graph.
@@ 7 Conclusions
Optimally distributing available resources and evaluating how, who, where, and when to intervene is crucial to mitigating climate change impacts. In this paper, we presented a novel data-driven approach for assessing the effectiveness of cash-based interventions in food emergencies through the lens of causal inference. We constructed a DAG to capture the dynamics of malnutrition under drought conditions and collected data characterizing the system. Our goal was to estimate the causal effects of cash-based interventions on reducing district-level global acute malnutrition across Somalia. Preliminary country-wise results did not reach statistical significance, although a singular district analysis did, prompting further steps: i) identifying more suitable treatment variables, ii) refining the causal graph with domain experts, iii) gaining insights on the spatio-temporal heterogeneity of impact of interventions through Conditional Average Treatment Effects (CATE) [54]. If data allows it, causal inference can be used to assess the efficacy of interventions in specific locations, supporting targeted aid where on-ground surveys are not feasible. The proposed approach could promote greater accountability and transparency amongst humanitarian actors, encouraging individuals to contribute to impactful and traceable aid.
@@ 8 Acknowledgments
This work was supported by the Fundación BBVA with the project ‘Causal inference in the human biosphere coupled system (SCALE)’, the Microsoft Climate Research Initiative through the Causal4Africa project, the European Union’s Horizon Europe Research and Innovation Program through the ThinkingEarth project (under Grant Agreement number 101130544) and the GVA PROMETEO AI4CS project on ‘AI for complex systems’ (2022-2026) with CIPROM/2021/056.